Problem 1.6

Two cars travel at the same constant speed. The first car travels for 10 minutes, the second for 27 minutes. If the second car travels 18 miles farther than the first, how far does each car travel? Solve algebraically.

Problem L.7

Two cars start from the same place at the same time. After two hours, one car has traveled 30Vo farther than the other. If one car's speed is 20 mph larger than the other car's speed, what are the speeds of the cars? Solve algebraically.
Answer #1

Problem 1.6:

Let suppose speed of both car is v mile/hr

First car travels for 10 min with speed v ,so it will cover a distance of

S1 =v×(10/60) miles

Second car travels for 27 min with velocity v so total distance covered by it

S2 =v×(27/60) miles

Now it Is given that

S2=S1+18

v×(27/60)=v×(10/60)+18

v×(17/60)=18

v=(18×60)/17=63.529 mile/hr

So now S1=63.529×(10/60)=10.588 miles

And S2=63.529×(27/60)=28.588 miles. Ans